OpenStudy (anonymous):

Solve the following system of equations. 3x +1/4y =-3/4 -1/2x+1/3y=8

OpenStudy (anonymous):

@TorilovesVampires

Vocaloid (vocaloid):

are you sure you typed out those equations correctly? they look somewhat odd to me

Vocaloid (vocaloid):

ok, much better. we can start by multiplying the second equation by 6

OpenStudy (anonymous):

-3x+2y=48

Vocaloid (vocaloid):

good, now we have: 3x + (1/4)y = (-3/4) -3x + 2y = 48 now add the two equations together

OpenStudy (anonymous):

x+2.25y=47.25

OpenStudy (anonymous):

don't think I'm supposed to have decimals

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

almost 3x - 3x = ?

OpenStudy (shadowlegendx):

I know this one :)

OpenStudy (anonymous):

x

Vocaloid (vocaloid):

nope, not quite hint: 3 - 3 = 0 3x - 3x = ?

OpenStudy (anonymous):

0x

OpenStudy (shadowlegendx):

There is no such thing as 0x That would be just 0...

Vocaloid (vocaloid):

good, and 0x = 0 because 0 times anything is 0 so now we have: 2.25y = 47.25 y = ?

OpenStudy (anonymous):

21

OpenStudy (shadowlegendx):

2.25y is saying 2.25 times y. Negate that by dividing. Divide both sides by 2.25

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

y= 21

OpenStudy (anonymous):

what about solving for X

Vocaloid (vocaloid):

ok, now we know that y = 21 we can solve for x by plugging in y = 21 back into one of the other equations

OpenStudy (anonymous):

for both equations

Vocaloid (vocaloid):

-(1/2)x + (1/3)y = 8 -(1/2)x + (1/3)21 = 8 -(1/2)x + 7 = 8 x = ?

Vocaloid (vocaloid):

we only need to plug in y = 21 into one of the equations, since both will give us the same x value

OpenStudy (anonymous):

.5

OpenStudy (anonymous):

1/2

Vocaloid (vocaloid):

nope, not quite (-1/2)x + 7 = 8 subtract 7 from each side to get (-1/2)x = 1 x = ?

OpenStudy (anonymous):

-.5

Vocaloid (vocaloid):

nope, check your math again

Vocaloid (vocaloid):

1/(-1/2) = ?

OpenStudy (anonymous):

-2

Vocaloid (vocaloid):

good, so x = -2 and y = 21

OpenStudy (anonymous):

thanks!